OpenStudy (anonymous):

evaluate line integral of (x+y) ds where C is the straight-line segment x=t, y=(1-t), z=0, from (0,1,0) to (1,0,0)

OpenStudy (loser66):

From (0, 1,0) to (1, 0,0) . That is when t =0 , x = 0, y =1, z =0; when t =1 , x =1, y =0, z=0 Hence, \(0\leq t\leq 1\) Now the integral: (x +y) w.r.t t is x+y = t +1-t =1 Jacobian: \(\sqrt{(dx/dt)^2+(dy/dt)^2+(dz/dt)^2}dt \) \(= \sqrt{(1)^2+(1)^2+(0)^2}dt =\sqrt{2}dt\) Combine all \[\int_0^1 \sqrt{2}dt\] I think you can handle it from here.

OpenStudy (anonymous):

thanks